OpenStudy (anonymous):

Write the equation of the line that passes through (1, -3) and is perpendicular to the line 2x + 3y = 10

OpenStudy (nikato):

WELCOME TO OPENSTUDY!

OpenStudy (nikato):

can you find the slope in 2x+3y=10

OpenStudy (anonymous):

-2x/3

OpenStudy (nikato):

ok. so slope is -2/3 what do you know about the slope of the line perpendicular to it?

OpenStudy (anonymous):

3/2

OpenStudy (nikato):

Yes. And you can plug that into the point-slope form

OpenStudy (nikato):

y-y1= m(x-x1) Where m=3/2 And (x1,y1)= (1,-3)

OpenStudy (anonymous):

thanks

OpenStudy (nikato):

You're welcome!